4. How many milliliters of a o 250 solution can be prepared by dissolving g of NaCI in water? 400 M many grams of lithium bromide LiBr could be recovered by evaporating 550. mL of 20.0 percent LiBr solution to dryness (d 1.34 g/mL)? 6. How many milliliters of M Hol is needed to prepare 500. mL of a 0.150 M HCI solution? 7. A sample of potassium hydrogen phthalate, HKCsH4o. weighing 0.512 g was dissolved in water and titrated with 24.82 mL of an NaoH solution. Calculate the molarity of the NaoH solution. 8. How many grams of HNO, are in 75. mL of concentrated (18 M) HNo, solution?
Answer #1

mol. wt of NaCl = 58.44g/mol

no of moles = 4/58.44 = 0.068

molarity = moles of solute/ litre of solution

litre of solution = 0.068/0.25 = 0.2737 litre

or 273.7 mL of water is required.